OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x))=x and g(f(x))=x f(x)=x-7/x+3 and g(x)= -3x -7/x-1

OpenStudy (jdoe0001):

so, are you confused where?

OpenStudy (anonymous):

Honestly the whole thing. I don't know why I am confused on this one but I am completely stuck

OpenStudy (jdoe0001):

well, f(x) = 2x+1 g(x) = WHATEVER f( g(x) ) = 2(WHATEVER) +1

OpenStudy (anonymous):

Please break it down for me for me

OpenStudy (jdoe0001):

whenever is said that say, f( g(x) ) f() is really CONTAINING g(x), or one can say, f(x) is HOSTING the g(x) GUEST

OpenStudy (jdoe0001):

so, every "x" or variable, in the HOST function, will be replaced by the GUEST function

OpenStudy (anonymous):

So how can I put this into the question I have? please help me work this question out

OpenStudy (jdoe0001):

do you know how to add fractions?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

$$\bf f(x)=\cfrac{x-7}{x+3}\\ g(x)=\color{red}{\cfrac{-3x -7}{x-1}}\\ f(g(x)) = \cfrac{\left(\color{red}{\cfrac{-3x -7}{x-1}}\right)-7}{\left(\color{red}{\cfrac{-3x -7}{x-1}}\right)+3} $$

OpenStudy (jdoe0001):

if g(x) is the inverse of f(x) those fractions addition will give you "x" only

OpenStudy (anonymous):

So I wrote it down but what would you get?

OpenStudy (jdoe0001):

$$f(g(x)) = \cfrac{\left(\cfrac{-3x -7}{x-1}\right)-7}{\left(\cfrac{-3x -7}{x-1}\right)+3}\\ \large \bf {\cfrac{ \frac{-3x-7-7x+7}{x-1}}{ \frac{-3x-7+3x-3}{x-1}} } $$

OpenStudy (jdoe0001):

what would that give you?

OpenStudy (anonymous):

-10x/x-1 -10/x-1

OpenStudy (jdoe0001):

$$\bf \cfrac{ \frac{-10x}{x-1}}{ \frac{-10}{x-1}} \implies \frac{-10x}{x-1} \times \frac{x-1}{-10} $$

OpenStudy (jdoe0001):

so, there :), is the inverse

OpenStudy (jdoe0001):

and you do the same with the g( f(x) )

OpenStudy (anonymous):

just x?

OpenStudy (jdoe0001):

yes, if f( g(x) ) = x, then g(x) is the inverse, or \(\bf f^{-1}(x) \) of f(x)

OpenStudy (anonymous):

could you please help me with the other one also?

OpenStudy (jdoe0001):

recall that => \(\bf \huge \cfrac{ \frac{a}{b}}{\frac{c}{d}} \implies \frac{a}{b} \times \frac{d}{c}\)

OpenStudy (anonymous):

so for the other one instead there be a x-1 in the denominator there would be x+3?

OpenStudy (jdoe0001):

the other one is g( f(x) ) so the HOST function is g(x), and the GUEST function is f(x) so any "x"'s in g(x) will be replaced for whatever f(x) is

OpenStudy (jdoe0001):

$$\bf f(x)=\frac{x-7}{x+3}\\ g(x)=\frac{-3x -7}{x-1}\\ g(f(x)) = \large \cfrac{-3\left(\frac{x-7}{x+3}\right) -7}{\left(\frac{x-7}{x+3}\right)-1} $$

OpenStudy (anonymous):

please can you tel me the answer. I am beyond confused

OpenStudy (anonymous):

so it would be x again?

OpenStudy (jdoe0001):

the exercises in themselves are meaningless, and so is the answer what makes them useful is solving them I mean, I can get the answer quickly, but that's not going to help you

OpenStudy (jdoe0001):

the idea behind them is that you understand what happens and why and where

OpenStudy (jdoe0001):

I mean, if all your teacher wanted was the answers filled in, she would have handed out to all students the answer sheet too, and that'd be all

OpenStudy (anonymous):

I understand but please just this time I've never been this confused ever I usually get it but this one I just don't

OpenStudy (jdoe0001):

well, is really just one function added to another at the "x" so really all "x" on the HOST are just PLACEHOLDERS for the GUEST function

OpenStudy (jdoe0001):

so if say, f(x) = x + 3x + 1 g(x) = WHATEVER f( g(x) ) = (WHATEVER)+3(WHATEVER) + 1

OpenStudy (jdoe0001):

it just happens that in this one, both functions are rationals, or fractions so you end up with a fraction over a fraction

OpenStudy (jdoe0001):

so, you add the fraction atop by itself then you add the one below by itself then just flip the denominator as \(\bf \large \cfrac{ \frac{a}{b}}{\frac{c}{d}} \implies \frac{a}{b} \times \frac{d}{c}\)

OpenStudy (jdoe0001):

which is really all I'm doing

OpenStudy (jdoe0001):

$$\bf \huge { g(f(x)) = \cfrac{-3\left(\frac{x-7}{x+3}\right) -7}{\left(\frac{x-7}{x+3}\right)-1}\\ \implies \cfrac{ \frac{-3x+21-7x-21}{x+3}}{ \frac{x-7-x-3}{x+3}} } $$

OpenStudy (anonymous):

what would you have to do after that?

OpenStudy (jdoe0001):

well, notice atop, is a fraction below, a fraction so you flip the denominator upside-down and multiply by the numerator, like \(\bf \large \cfrac{ \frac{a}{b}}{\frac{c}{d}} \implies \frac{a}{b} \times \frac{d}{c}\)

OpenStudy (anonymous):

-3x+21-7x-21/x+3 X x+3/x-7-x-3

OpenStudy (jdoe0001):

yes, the x+3 will cancel out with each other

OpenStudy (anonymous):

-3x-3+7x-3 so how do I connect with the x from before?

OpenStudy (jdoe0001):

well, is -3x-3-7x-3

OpenStudy (jdoe0001):

hmm wait

OpenStudy (anonymous):

oh my bad. So is that the final answer?

OpenStudy (jdoe0001):

the expression above is -3x+21-7x-21 => -3x-7x+21-21 => -10x the expression below is x-7-x-3 => x-x-7-3 => -10

OpenStudy (jdoe0001):

$$\bf \large\cfrac{ \frac{-3x+21-7x-21}{x+3}}{ \frac{x-7-x-3}{x+3}} \implies \frac{-3x+21-7x-21}{\cancel{x+3}} \times \frac{\cancel{x+3}}{x-7-x-3}\\ \frac{-10x}{-10} $$

OpenStudy (anonymous):

which is x

OpenStudy (jdoe0001):

yes, which means g( f(x) ) = x thus g(x) is the inverse of f(x)

OpenStudy (anonymous):

Got the final answer thank you so much! :)